The spot rate of euro against US$ is $1.4250/€. The rate of inflation in the U.S. and Euro area are expected to be 3% and 2.5% respectively. Based on the theory of Relative PPP, the expected spot exchange rate of euro against the US$ for next year is ____________.
Answer #1

Purchasing Power Parity Theory is given by

1+Ih / 1+If = F1/S0

where

Ih - inflation rate in home country =3%

If - inflation rate in foreign country = 2.50%

F1 - 1 year forward rate =?

S0 - Spot RAte =1.4250

Here exchange rate is given in the format of  $/€. Hence, $(first currency) is the home country and €(second currency) is the foreign country.

1.03/1.025 = F1/1.425

F1/1.425 = 1.00487804878

F1 = 1.00487804878*1.425

= 1.4320

PPT states that high Inflation rate in a country is offset by depreciation in the currency of that country. So here $ should depreciate.

Here forward rate>spot rate. That is product in the exchange rate, € is appreciating and $ is depreciating, which is in line with PPT.